OpenStudy (anonymous):

Solve the inequality. 3 – 2a < 7 A. a < –2 B. a < 2 C. a > –2 D a > 2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Mesopretty can you tell me how you got that so i can do it next time

OpenStudy (anonymous):

Simplifying 3 + -2a = 7 Solving 3 + -2a = 7 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -2a = 7 + -3 Combine like terms: 3 + -3 = 0 0 + -2a = 7 + -3 -2a = 7 + -3 Combine like terms: 7 + -3 = 4 -2a = 4 Divide each side by '-2'. a = -2 Simplifying a = -2 See similar equations: | x-9=25 | | 126-4x=4x+14 | | 4w=20 | | -3=4x+10 | | 16x-3=11x+12 | | 4x+2y=8 | | 5t-8-2t+3= | | 3z+1=z-6 | | 6x-3y=5 | | 9(3x-5)=8(35-x)+25 | | -104-8x=8x+24 | | y-3=7 | | 17y-7=83-y | | 30+X^2+4x=0 | | 8X+7Y-3X+4-Y= | | X^2=-30+-4x | | 4-5p-2-3p= | | 2(p+4)+1=5p | | 5(x-3)-25=0 | | 2x+2y=180 | | a-4b+3a= | | -3(4n+8)=3 | | 7(x)+9=15 | | -4(5n+1)=12 | | -5x-3y+4z=5 | | 16x(2)-25=0 | | 5x+3=-(x-27) | | 8x+8=48 | | 7+4(x-6)=6-7(x-5) | | -4(5n+2)=4 | | 6x-10x+2=-2x+6 | | 2+9(x-4)=3-9(x-3) | | 7*5x+9+(x+9)=12 | | x^2-7=0 | | 2x+11=9x-3 | | 15m=8 | | -14p-3*(3-5p)=3(p-6)-3 | | 5a-c=43 | | -4x+2=2 | | 5n-8=4+8n | | -4x+7=x+2 | | 5x+3y=7 | | 2x+9-5x+3=45 | | z-x=y | | 3p-2*1p=6*-4p-30 | | a+b=c | | 7.6x-2.7=12.5 | | 14=0.1P | | 14-9x=178-37x | | 3T+18-7=41 | | 5s+t-2s+6= | | -3(2x-4)=24 | | 14m-10=3(4+3) | | 4p-q+r+2q-r= | | 3(-2-3x)=9x-4 | | 45-6x=3x+19 | | 5(2z-1)-4(z+7)=2(z+1) | | 6=60-21x | | 11x-29=92 | | 136-5x=31x+20 | | 2x+y-z+3x-2y= | | 4x+32=6x+12 | | 3q+7q+9q= | | 10x-50=6x-6 | | 4m=11 | Equations solver categories Equations solver - equations involving one unknown Quadratic equations solver Percentage Calculator - Step by step Derivative calculator - step by step Greatest Common Factor Least Common Multiple System of equations - step by step solver Simplifying 3 + -2a = 7 Solving 3 + -2a = 7 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -2a = 7 + -3 Combine like terms: 3 + -3 = 0 0 + -2a = 7 + -3 -2a = 7 + -3 Combine like terms: 7 + -3 = 4 -2a = 4 Divide each side by '-2'. a = -2 Simplifying a = -2 See similar equations: | x-9=25 | | 126-4x=4x+14 | | 4w=20 | | -3=4x+10 | | 16x-3=11x+12 | | 4x+2y=8 | | 5t-8-2t+3= | | 3z+1=z-6 | | 6x-3y=5 | | 9(3x-5)=8(35-x)+25 | | -104-8x=8x+24 | | y-3=7 | | 17y-7=83-y | | 30+X^2+4x=0 | | 8X+7Y-3X+4-Y= | | X^2=-30+-4x | | 4-5p-2-3p= | | 2(p+4)+1=5p | | 5(x-3)-25=0 | | 2x+2y=180 | | a-4b+3a= | | -3(4n+8)=3 | | 7(x)+9=15 | | -4(5n+1)=12 | | -5x-3y+4z=5 | | 16x(2)-25=0 | | 5x+3=-(x-27) | | 8x+8=48 | | 7+4(x-6)=6-7(x-5) | | -4(5n+2)=4 | | 6x-10x+2=-2x+6 | | 2+9(x-4)=3-9(x-3) | | 7*5x+9+(x+9)=12 | | x^2-7=0 | | 2x+11=9x-3 | | 15m=8 | | -14p-3*(3-5p)=3(p-6)-3 | | 5a-c=43 | | -4x+2=2 | | 5n-8=4+8n | | -4x+7=x+2 | | 5x+3y=7 | | 2x+9-5x+3=45 | | z-x=y | | 3p-2*1p=6*-4p-30 | | a+b=c | | 7.6x-2.7=12.5 | | 14=0.1P | | 14-9x=178-37x | | 3T+18-7=41 | | 5s+t-2s+6= | | -3(2x-4)=24 | | 14m-10=3(4+3) | | 4p-q+r+2q-r= | | 3(-2-3x)=9x-4 | | 45-6x=3x+19 | | 5(2z-1)-4(z+7)=2(z+1) | | 6=60-21x | | 11x-29=92 | | 136-5x=31x+20 | | 2x+y-z+3x-2y= | | 4x+32=6x+12 | | 3q+7q+9q= | | 10x-50=6x-6 | | 4m=11 | Equations solver categories Equations solver - equations involving one unknown Quadratic equations solver Percentage Calculator - Step by step Derivative calculator - step by step Greatest Common Factor Least Common Multiple System of equations - step by step solver

OpenStudy (anonymous):

thanks for all the information

OpenStudy (anonymous):

Welcome